An electron that has a velocity with x component 2.7 × 106 m/s and y component 2.3 × 106 m/s moves through a uniform magnetic field with x component 0.033 T and y component -0.21 T. (a) Find the magnitude of the magnetic force on the electron. (b) Repeat your calculation for a proton having the same velocity.
Answer #1

A.

F = q*VXB

VXB = \begin{bmatrix} i & j & k\\ 2.7*10^6& 2.3*10^6 & 0\\ 0.033& -0.21 &0 \end{bmatrix}

VXB = (-2.7*0.21*10^6 - 0.033*2.3*10^6) k

VXB = -6.429*10^5 k

Now for electrons

F = q*VXB = (-1.6*10^-19)*(-6.429*10^5) = 1.029*10^-13 N

|F| = 1.029*10^-13 N

B.

there will be no change in to magnitude of magnetic force, since magnitude of charge is same for both proton and electron.

F = -1.029*10^-13 N

|F| = 1.029*10^-13 N